OpenStudy (anonymous):

find the integers,given the following sum is -12 product is 35

OpenStudy (anonymous):

-7 and -5

OpenStudy (chaise):

x+y=-12 xy=35 -7 and -5

OpenStudy (dumbcow):

i agree

OpenStudy (anonymous):

x+y=-12 x-y=\[\sqrt{(x+y)^{2}-4xy}\]=\[\sqrt{(-12)^{2}-4*35}=2\] solve these 2 equaions and you will get x and y as -7 and -5